OpenStudy (anonymous):

Prove that If B is invertible, then det(B^-1*A*B)=det(A). Thanks for any help :)

OpenStudy (kainui):

Do you know any rules of determinants?

OpenStudy (kainui):

how about if you have two matrices G and H then: det(GH)=det(G)*det(H)

OpenStudy (anonymous):

Is it equal to the identity matrix?

OpenStudy (anonymous):

I mean...det(B^-1AB)=det(B-1)det(A)det(B)=Det(A)*I=det(A)

OpenStudy (anonymous):

is det(B-1)*det(B)=I?

OpenStudy (anonymous):

@Kainui

OpenStudy (kainui):

Well luckily we can show this quite easily! Check this out! \[\LARGE 1=\det(I)=\det(BB^{-1})=\det(B)\det(B^{-1})\]

OpenStudy (anonymous):

YAY!! Thankyou :)